# DFML: Decentralized Federated Mutual Learning

### v1

@@ 5.2 Homogeneous Architectures
5.3 Heterogeneous Architectures with Restrictions
In this set of experiments, some restrictions are applied to heterogeneous architectures. Table 3 demonstrates that DFML consistently outperforms all baselines in terms of global accuracy across three different architectures, two datasets, and both IID and non-IID data distributions. Another observation is that the partial training baselines outperformed decentralized FedAvg, which is expected as more knowledge is shared by averaging overlapping parameters of the models rather than only averaging models with the same architectures. Furthermore, we implemented decentralized Federated Dropout; however, the results are not reported as it did not work. The poor performance is attributed to local models being assigned random parameters from the global model generated at the aggregator.
5.4 Heterogeneous Architectures without Restrictions
After confirming that our proposed DFML competes with the prevalent baselines in DFL, we showcase the strength of DFML in knowledge transfer using nonrestrictive heterogeneous models. Additionally, we evaluate DFML under different $N$ clients. From Table 4, it is evident that across various $N$ clients, datasets, and data distributions, DFML consistently achieves superior global accuracy compared to the baseline. Table 5 shows that DFML requires fewer communication rounds to reach the same accuracy as decentralized FedAvg does at specific rounds. The communication cost per round for all experiments is $50$, involving the transmission of $25$ models to and from the aggregator. Appendix D provides further experimental analysis on DFML. A visual illustration of the convergence speedup achieved by DFML compared to the baselines is presented in Appendix D.1. Moreover, a comparison between DFML and a decentralized version of FML is provided in Appendix D.5.
@@ 5.4 Heterogeneous Architectures without Restrictions
Method CE / WSM Cyclical $\boldsymbol{\alpha}$ IID non-IID IID non-IID IID non-IID
@@ 6 Analysis
In this section, we first demonstrate the effect of using different fixed $\alpha$ values versus cyclically adjusting it. Second, we present the impact of cyclical $\alpha$ on global accuracy by evaluating the regular models. Subsequently, we examine the behavior exhibited by both regular and peak models as $\alpha$ changes, and highlight the ability of the peak models to capture the peaks in global performance. Last, we conduct an analysis to understand the effects of different supervision signals and cyclical $\alpha$ on global accuracy.
Figure 5: Performance comparison of the regular models against cyclically oscillating $\alpha$ with half cycles. The supervision signal used is CE. Refer to Appendix E for full-cycle $\alpha$ oscillation.
@@ 6.1 Regular vs Peak Models
Figure 5 shows the fluctuations in global performance as $\alpha$ is cyclically varied. When $\alpha=0$, representing only the supervision signal in the objective function, global accuracy is at its lowest. This is attributed to models being optimized solely toward the aggregator’s local data. However, as $\alpha$ increases and eventually reaches its maximum defined value, models gain knowledge from each other through knowledge distillation, resulting in a peak in global accuracy. The decline in global accuracy starts when the distillation signal diminishes, and the supervisory signal takes over. To maintain global performance at the peaks, the peak models are used. Initially, the peak models are updated regularly with every regular model update until the first maximum $\alpha$ value is reached. After that, the peak models are only updated when the regular models are aggregated using the highest value of $\alpha$. Figure 6 depicts the global accuracy of both regular and peak models, highlighting the stability achieved by the peak models throughout training.
@@
6.2 Fixed vs Cyclical $\boldsymbol{\alpha}$
As shown in Figure 5, the highest global accuracy is consistently achieved when $\alpha$ reaches its maximum defined value. Now, we address the impact of using different fixed values for $\alpha$ and whether setting $\alpha=1$ throughout training yields similar performance as cyclically changing $\alpha$. Figure 7 demonstrates that using different fixed values of $\alpha$, under different distribution shifts, results in varied performance levels. Furthermore, fixing $\alpha=1$ leads to the worst global performance because when only the distillation signal is present throughout training without any supervision, noise will be propagated. This results in experts teaching each other incorrect information.
6.3 Supervision Signal and Cyclical $\boldsymbol{\alpha}$
The impact of different supervision signals (CE and WSM) and cyclical $\alpha$ on DFML is presented in Table 6. Results indicate that the use of WSM primarily enhances the global accuracy in non-IID data distribution shifts. Furthermore, the addition of cyclical $\alpha$ on top of CE and WSM further improves global accuracy. The best outcomes are mostly reported when both WSM and cyclical $\alpha$ are applied. Additional experimental analysis on cyclical knowledge distillation can be found in Appendix E.
@@ 7 Conclusion
We proposed DFML, a framework that supports a decentralized knowledge transfer among heterogeneous models, without architectural constraints or reliance on additional data. DFML outperformed state-of-the-art baselines in addressing model and data heterogeneity in DFL, showcasing better convergence speed and global accuracy.
Appendix A Dataset Distributions
@@
Appendix B Implementation Details
In our setup, we assume a star topology where all clients can send and receive models from any other client. Each communication round involves randomly selecting 50% of the clients as senders $S$, with an additional client randomly chosen as the aggregator $a$ (unless specified otherwise). We utilize SGD optimizer for each client with momentum 0.9 and weight decay 5e-4. The learning rate is selected from $\{0.1,0.01,0.001\}$. The batch size is set to 64. For the cyclic $\alpha$ scheduler, we apply cosine annealing. The initial oscillating period is set to 10 and is incrementally increased after each completion. $\alpha$ is oscillated from 0 to a maximum value selected between $\{0.8,0.9,1.0\}$. Figure 9 illustrates an example of the behavior of $\alpha$ throughout training. The number of mutual learning epochs $K$, performed at the aggregator, is set to 10. Moreover, the temperature is configured to 1. All experiments are repeated for 3 trials with random seeds. With the existence of an $\alpha$ and period schedulers, aggregators need to be aware of communication round $t$, and the round when the period was last updated to compute the period and $\alpha$ value. To achieve this, each aggregator needs to communicate these two values to the next aggregator.
@@ Appendix B Implementation Details
Appendix C Baselines
C.1 Decentralized FedAvg
@@
Algorithm 2 Decentralized FedAvg
@@
           $W_{n}\leftarrow$ locally train $W_{n}$ using Equation (2)
@@
        $W_{g}^{u}\leftarrow$ Aggregate homogeneous models, for all $W_{n}\in u$, according to Equation (4)
@@
C.2 Decentralized Partial Training
@@
C.2.1 Decentralized Federated Dropout
In each communication round $t$ within Federated Dropout (Caldas et al., 2018), sub-models (local models) are extracted from the centralized $W_{g}$ based on a random selection process. The parameters representing the sub-models that are randomly chosen through this selection are then transmitted to clients for local training. After local training, the updated parameters are sent back to $W_{g}$ for aggregation. The random extraction scheme for layer $i$ of sub-model $W_{n}$ for client $n$ is extracted from $W_{g}$ as follows:
@@ C.2.1 Decentralized Federated Dropout
Our derived decentralized version of Federated Dropout involves generating random indices (Equation (6)) at the aggregator, guided by the largest available model ($W_{g}$). Subsequently, each model is assigned a random set of indices equivalent to its size. The aggregation process is then carried out using these randomly selected indices. Once the aggregation is finalized, sub-models are created from $W_{g}$ using the same set of indices. Finally, these sub-models are transmitted back to the respective participating clients.
C.2.2 Decentralized HeteroFL
Unlike Federated Dropout, HeteroFL (Diao et al., 2020) consistently extracts sub-models from a predefined section of $W_{g}$. Specifically, HeteroFL extracts sub-models from $W_{g}$ starting from index 0 up to the maximum layer size of $W_{n}$. The extraction scheme is defined as follows:
@@ C.2.2 Decentralized HeteroFL
C.2.3 Decentralized FedRolex
In FedRolex (Alam et al., 2022), local clients are initially generated from the global model beginning from index 0 and extending up to the local models’ capacity. In the first communication round, the sub-models are generated similarly to HeteroFL. However, in each subsequent communication round, the starting point of the indices shifts to the right. The sub-model extraction in FedRolex is defined as follows:
@@ C.2.3 Decentralized FedRolex
In decentralized FedRolex, the size of $W_{g}$ is determined by the largest available model, and as a result, the rightward shift in indices is computed based on the current communication round $t$ and $J_{i}$ of the selected $W_{g}$. The indices are calculated using $\mathcal{X}^{t}_{n,i}$ from Equation (8). These indices are utilized for aggregation and to extract the updated local models after aggregation. In decentralized FedRolex, since aggregators must be aware of $t$ to compute the indices, each aggregator needs to communicate $t$ to the next aggregator.
@@
           $W_{n}\leftarrow$ locally train $W_{n}$ using Equation (2)
@@
     Local models are assigned indices $\mathcal{X}_{n,i}$ for all $i$ and $n\in\mathcal{P}$, where $\mathcal{X}_{n,i}$ is from Equation (6) or Equation (7) or Equation (8)
@@ C.3 Decentralized FML
Appendix D DFML: Further Analysis
D.1 Convergence Speedup
We illustrate in Figures 10, 11, and 12 the convergence speedup achieved by DFML compared to the baselines under three heterogeneity settings: homogeneous architectures, restrictive heterogeneous architectures, and nonrestrictive heterogeneous architectures; respectively.

### v2

@@ 4.4.2 Homogeneous Architectures
4.4.3 Heterogeneous Architectures with Restrictions
Table 5: Global accuracy comparison using restrictive heterogeneous architectures. The experiments are conducted using ResNet architectures with $50$ clients and $25$ senders.
FMNIST Caltech101 Oxford Pets Stanford Cars
Method IID non-IID IID non-IID IID non-IID IID non-IID
Dec. FedAvg 89.07$\pm$ 0.05 73.53$\pm$ 0.13 36.74$\pm$ 0.61 22.83$\pm$ 0.34 10.25$\pm$ 0.43 9.21$\pm$ 0.31 2.83$\pm$ 0.24 3.44$\pm$ 0.21
Dec. HeteroFL 91.06$\pm$ 0.15 85.49$\pm$ 0.10 49.05$\pm$ 0.21 38.04$\pm$ 0.16 23.35$\pm$ 0.14 17.68$\pm$ 0.12 5.55$\pm$ 0.14 6.25$\pm$ 0.10
Dec. FedRolex 91.26$\pm$ 0.20 86.06$\pm$ 0.19 49.14$\pm$ 0.32 37.75$\pm$ 0.22 22.67$\pm$ 0.14 18.39$\pm$ 0.15 5.88$\pm$ 0.11 6.09$\pm$ 0.09
DFML (Ours) 92.01$\pm$ 0.02 87.31$\pm$ 0.03 60.05$\pm$ 0.04 50.07$\pm$ 0.06 40.42$\pm$ 0.16 31.25$\pm$ 0.11 41.12$\pm$ 0.10 25.31$\pm$ 0.09
In this set of experiments, some restrictions are applied to heterogeneous architectures. Tables 4 and 5 demonstrate that DFML consistently outperforms all baselines in terms of global accuracy across different architectures, datasets, and both IID and non-IID data distributions. Another observation is that the partial training baselines outperformed decentralized FedAvg, which is expected as more knowledge is shared by averaging overlapping parameters of the models rather than only averaging models with the same architectures. Furthermore, we implemented decentralized Federated Dropout; however, the results are not reported as it did not work. The poor performance is attributed to local models being assigned random parameters from the global model generated at the aggregator.
4.4.4 Heterogeneous Architectures without Restrictions
After confirming that our proposed DFML competes with the prevalent baselines in DFL, we showcase the strength of DFML in knowledge transfer using nonrestrictive heterogeneous models. Additionally, we evaluate DFML under different $N$ clients. From Tables 6 and 7, it is evident that across various $N$ clients, datasets, and data distributions, DFML consistently achieves superior global accuracy compared to the baseline. Table 8 shows that DFML requires fewer communication rounds to reach the same accuracy as decentralized FedAvg does at specific rounds. The communication cost per round for all experiments is $50$, involving the transmission of $25$ models to and from the aggregator. Appendix A.4 provides further experimental analysis on DFML. A visual illustration of the convergence speedup achieved by DFML compared to the baselines is presented in Appendix A.4.4. Moreover, a comparison between DFML and a decentralized version of FML is provided in Appendix A.4.11.
@@ 4.4.4 Heterogeneous Architectures without Restrictions
Table 7: Global accuracy comparison using nonrestrictive heterogeneous architectures. The experiments are conducted using EfficientNet architectures with $10$ clients and $5$ senders.
Caltech101 Oxford Pets Stanford Cars
Method IID non-IID IID non-IID IID non-IID
Dec. FedAvg 54.75$\pm$ 0.24 30.92$\pm$ 0.16 18.34$\pm$ 0.31 15.34$\pm$ 0.21 6.07$\pm$ 0.25 11.83$\pm$ 0.12
DFML (Ours) 83.42$\pm$ 0.20 65.70$\pm$ 0.13 68.90$\pm$ 0.22 42.94$\pm$ 0.13 70.29$\pm$ 0.19 50.63$\pm$ 0.23
@@
Method CE / WSM Cyclical $\bm{\alpha}$ IID non-IID IID non-IID IID non-IID
@@
Figure 5: Performance comparison between the different architecture clusters in both DFML and decentralized FedAvg. In this experiment, five nonrestrictive heterogeneous architectures are distributed among 50 clients. C0, C1, C2, C3, and C4 represent the global accuracy average of all models with CNN architectures [32, 64, 128, 256], [32, 64, 128], [32, 64], [16, 32, 64], and [8, 16, 32, 64], respectively.
4.4.5 Performance per Cluster of Architectures
Figure 5 presents the performance of each architecture cluster in DFML compared to decentralized FedAvg. In this experiment, five different CNN architectures are distributed among 50 clients. The figure demonstrates that the bigger the architecture size, the higher the attained global accuracy. Moreover, all clusters in DFML surpasses their corresponding clusters in decentralized FedAvg.
4.4.6 High Model Heterogeneity
To showcase the scalability of DFML with model heterogeneity, we conduct an experiment involving 50 clients with significant model heterogeneity. We compare the results obtained by DFML to decentralized FedAvg. In this experiment, ten different architectures are deployed and selected from Table 10: the two largest ViT architectures, the four largest ResNet architectures, and the four largest CNN architectures under the $H2$ category. Figure 6 demonstrates that DFML performs effectively under heavy model heterogeneity conditions and greatly outperforms decentralized FedAvg.
Figure 6: Comparison between DFML and decentralized FedAvg under significant model heterogeneity. Ten different architectures are distributed among 50 clients including: 2 ViT, 4 ResNet, and 4 CNN architectures.
@@ 5 Analysis
In this section, we first demonstrate the effect of using different fixed $\alpha$ values versus cyclically adjusting it. Second, we present the impact of cyclical $\alpha$ on global accuracy by evaluating the regular models. Subsequently, we examine the behavior exhibited by both regular and peak models as $\alpha$ changes, and highlight the ability of the peak models to capture the peaks in global accuracy. Last, we conduct an analysis to understand the effects of different supervision signals and cyclical $\alpha$ on global accuracy.
Figure 7: Performance comparison of the regular models against cyclically oscillating $\alpha$ with full and half cycles, respectively. The supervision signal used is CE.
@@ 5.1 Regular vs Peak Models
Figure 7 shows the fluctuations in global accuracy as $\alpha$ is cyclically varied. When $\alpha=0$, representing only the supervision signal in the objective function, global accuracy is at its lowest. This is attributed to models being optimized solely toward the aggregator’s local data. However, as $\alpha$ increases and eventually reaches its maximum defined value, models gain knowledge from each other through knowledge distillation, resulting in a peak in global accuracy. The decline in global accuracy starts when the distillation signal diminishes, and the supervisory signal takes over. To maintain global accuracy at the peaks, the peak models are used. Initially, the peak models are updated regularly with every regular model update until the first maximum $\alpha$ value is reached. After that, the peak models are only updated when the regular models are aggregated using the highest value of $\alpha$. Figure 8 depicts the global accuracy of both regular and peak models, highlighting the stability achieved by the peak models throughout training.
5.2 Fixed vs Cyclical $\alpha$
As shown in Figure 7, the highest global accuracy is consistently achieved when $\alpha$ reaches its maximum defined value. Now, we address the impact of using different fixed values for $\alpha$ and whether setting $\alpha=1$ throughout training yields similar performance as cyclically changing $\alpha$. Figure 9 demonstrates that using different fixed values of $\alpha$, under different distribution shifts, results in varied performance levels. Furthermore, fixing $\alpha=1$ leads to the worst global accuracy because when only the distillation signal is present throughout training without any supervision, noise will be propagated. This results in experts teaching each other incorrect information.
@@ 5.2 Fixed vs Cyclical $\alpha$
5.3 Supervision Signal and Cyclical $\alpha$
The impact of different supervision signals (CE and WSM) and cyclical $\alpha$ on DFML is presented in Table 9. Results indicate that the use of WSM primarily enhances the global accuracy in non-IID data distribution shifts. Furthermore, the addition of cyclical $\alpha$ on top of CE and WSM further improves global accuracy. The best outcomes are mostly reported when both WSM and cyclical $\alpha$ are applied. Additional experimental analysis on cyclical knowledge distillation can be found in Appendix A.5.
6 Limitations
While DFML presents significant advancements over existing baselines, it does have certain limitations that need to be acknowledged. One notable limitation is its computational expense. DFML requires aggregators to possess sufficient memory to receive multiple models and enough computational power to perform mutual learning, involving both forward and backward passes on all models at the aggregator. Another limitation is the lack of convergence analysis within in this paper. While proving the convergence properties of DFML is beyond the current scope, proving the convergence properties of DFML is left for future work.
@@ 7 Conclusion
We proposed DFML, a framework that supports a decentralized knowledge transfer among heterogeneous models, without architectural constraints or reliance on additional data. DFML overcomes common centralization issues such as communication bottlenecks and single points of failure, making it a robust alternative for real-world applications. DFML outperformed state-of-the-art baselines in addressing model and data heterogeneity in DFL, showcasing better convergence speed and global accuracy.
Appendix A Appendix
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 10: Data partitions based on IID and non-IID distribution with clients $N=\{10,50,100\}$ for both CIFAR-10 and CIFAR-100 datasets. The size of the red circle represents the magnitude of data samples for each class label in each client.
A.1 Dataset Distributions
@@
A.2 Implementation Details
In our setup, we assume a star topology where all clients can send and receive models from any other client. Each communication round involves randomly selecting 50% of the clients as senders $S$, with an additional client randomly chosen as the aggregator $a$ (unless specified otherwise). We utilize SGD optimizer for each client with momentum 0.9 and weight decay 5e-4. The learning rate is selected from $\{0.1,0.01,0.001\}$. The batch size is set to 8 for the EfficientNet experiments, 16 for the ResNet experiments using Caltech101, Oxford Pets, and StanfordCars datasets, and batch size of 64 is used for all other experiments,. For the cyclic $\alpha$ scheduler, we apply cosine annealing. The initial oscillating period is set to 10 and is incrementally increased after each completion. $\alpha$ is oscillated from 0 to a maximum value selected between $\{0.8,0.9,1.0\}$. Figure 11 illustrates an example of the behavior of $\alpha$ throughout training. The number of mutual learning epochs $K$, performed at the aggregator, is set to 10. Moreover, the temperature is configured to 1. All experiments are repeated for 3 trials with random seeds. With the existence of an $\alpha$ and period schedulers, aggregators need to be aware of communication round $t$, and the round when the period was last updated to compute the period and $\alpha$ value. To achieve this, each aggregator needs to communicate these two values to the next aggregator.
@@ A.2 Implementation Details
EfficientNet B0
B1
B2
B3
B4
A.3 Baselines
A.3.1 Decentralized FedAvg
@@
A.3.2 Decentralized FedProx
Decentralized FedProx is similar to decentralized FedAvg in that a subset of clients send their locally trained models to the aggregator for parameter averaging to form a global model, which is then sent back to the senders. However, FedProx adds a proximal term to the local training objective, which helps improve the method’s stability by restricting the local updates to be closer to the initial (global) model. The proximal term necessitates keeping a copy of the global model while performing local training. In decentralized setting with partial client participation, multiple versions of global models exist throughout the network. Decentralizing FedProx requires each client to treat its model as the global model before performing local training. Algorithm 2 provides a detailed explanation of decentralized FedProx. The hyperparameter $\mu$ is selected from $\{0.5,1,2\}$.
Algorithm 2 Decentralized FedAvg and FedProx
@@ A.3.2 Decentralized FedProx
           $W_{n}\leftarrow$ locally train $W_{n}$ using Equation 2
           $W_{n}\leftarrow$ locally train $W_{n}$ using Equation 2 + $\frac{\mu}{2}\left\|\widetilde{W}_{n}-W_{n}\right\|^{2}$
@@
        $W_{g}^{u}\leftarrow$ Aggregate homogeneous models, for all $W_{n}\in u$, according to Equation 5
@@
A.3.3 Decentralized Partial Training
@@
Decentralized Federated Dropout In each communication round $t$ within Federated Dropout (Caldas et al., 2018), sub-models (local models) are extracted from the centralized $W_{g}$ based on a random selection process. The parameters representing the sub-models that are randomly chosen through this selection are then transmitted to clients for local training. After local training, the updated parameters are sent back to $W_{g}$ for aggregation. The random extraction scheme for layer $i$ of sub-model $W_{n}$ for client $n$ is extracted from $W_{g}$ as follows:
@@
Our derived decentralized version of Federated Dropout involves generating random indices (Equation 7) at the aggregator, guided by the largest available model ($W_{g}$). Subsequently, each model is assigned a random set of indices equivalent to its size. The aggregation process is then carried out using these randomly selected indices. Once the aggregation is finalized, sub-models are created from $W_{g}$ using the same set of indices. Finally, these sub-models are transmitted back to the respective participating clients.
Decentralized HeteroFL Unlike Federated Dropout, HeteroFL (Diao et al., 2020) consistently extracts sub-models from a predefined section of $W_{g}$. Specifically, HeteroFL extracts sub-models from $W_{g}$ starting from index 0 up to the maximum layer size of $W_{n}$. The extraction scheme is defined as follows:
@@
Decentralized FedRolex In FedRolex (Alam et al., 2022), local clients are initially generated from the global model beginning from index 0 and extending up to the local models’ capacity. In the first communication round, the sub-models are generated similarly to HeteroFL. However, in each subsequent communication round, the starting point of the indices shifts to the right. The sub-model extraction in FedRolex is defined as follows:
@@
In decentralized FedRolex, the size of $W_{g}$ is determined by the largest available model, and as a result, the rightward shift in indices is computed based on the current communication round $t$ and $J_{i}$ of the selected $W_{g}$. The indices are calculated using $\mathcal{X}^{t}_{n,i}$ from Equation 9. These indices are utilized for aggregation and to extract the updated local models after aggregation. In decentralized FedRolex, since aggregators must be aware of $t$ to compute the indices, each aggregator needs to communicate $t$ to the next aggregator.
@@
           $W_{n}\leftarrow$ locally train $W_{n}$ using Equation 2
@@
     Local models are assigned indices $\mathcal{X}_{n,i}$ for all $i$ and $n\in\mathcal{P}$, where $\mathcal{X}_{n,i}$ is from Equation 7 or Equation 8 or Equation 9
@@ A.3.4 Decentralized FML
A.4 DFML: Further Analysis
A.4.1 Centralized vs Decentralized
We highlight a challenge of a serverless setting compared to having a centralized server in the network. For simplicity, we illustrate the difference using FedAvg. In centralized FedAvg, a global model exists at the server, and in each communication round, the global model is distributed to randomly selected clients. Even when partial client participation, there is always one version of the global model at the server. In a decentralized setting with partial participation, clients send their model to another client for aggregation and receive back the aggregated model, resulting in a version of the global model that differs from the previous round especially with different clients selected for participation. The existence of multiple versions of global models among clients in a serverless network affects the convergence speed compared to a centralized server where the latest version of the global model is transmitted to clients in every round. Figure 12 illustrates the convergence speed drop between a centralized and decentralized network using FedAvg and homogeneous architectures. Additionally, our proposed DFML challenges the centralized FedAvg. As mentioned in Section 1, there are other advantages of a serverless network compared to having a centralized server.
Figure 12: Performance comparison between centralized (with server) and decentralized (serverless) settings.
A.4.2 Fixed vs Random Aggregator
Figure 13 shows that using a fixed client to perform the aggregation significantly limits performance since the same local data is used in the knowledge transfer between models. In this experiment, the senders are randomly selected in each communication round. while the aggregator is always client 0.
Figure 13: Performance comparison between using the same client (client 0) and randomly selecting a client in each communication round to perform the aggregation.
A.4.3 Effect of Knowledge Distillation
We explore the effect of utilizing a distillation signal in the objective function compared to just having a supervision signal at the aggregator (Equation 1). Figure 14 illustrates the communication cost difference between having both supervision and distillation signals in the aggregator’s objective function versus just having the supervision signal. The figure shows that without the distillation signal, the computation cost is approximately an order of magnitude higher than with the addition of the distillation signal to the supervision signal at the aggregator, as described in Equation 1.
Figure 14: Performance comparison between using both distillation and supervision signals and only using the supervision signal at the aggregator to perform knowledge transfer.
A.4.4 Convergence Speedup
We illustrate in Figures 15, 16, and 17 the convergence speedup achieved by DFML compared to the baselines under three heterogeneity settings: homogeneous architectures, restrictive heterogeneous architectures, and nonrestrictive heterogeneous architectures; respectively.